OpenStudy (elisaneedshelp):

Hello Can somebody check an answer for me about finding consecutive integers please? :( :D

OpenStudy (elisaneedshelp):

Ok so i was given an explanation but it was wrong

OpenStudy (elisaneedshelp):

he meaning of consecutive means continuous {please don't be offended } so let the first term be 'y', the second one be 'y+1' and the third one be 'y+2' so according to the question, y+(y+1)+(y+2)=21 you don't need to worry as there is a '+' sign outside the bracket you just need to add = y+y+1+y+2=21 group the like terms [terms with same ending like 'y'] = y+y+y+1+2 =21 3y+3=21 {just to tell you y*y*y= y³ } now 3y=21-3 3y=18 y=18/3 y=6 so the integers are y=6, y+1=6+1=7 and y+2= 6+2=8 if you want to check=6+7=13+8=21 hence proved

OpenStudy (elisaneedshelp):

Somehow this is wrong and it was incorrect

OpenStudy (elisaneedshelp):

Three consecutive odd integers have a sum of 21. Find the integers.

OpenStudy (elisaneedshelp):

^thats the question

OpenStudy (elisaneedshelp):

the answer 6, 7, and 8 was wrong

OpenStudy (elisaneedshelp):

@escamer :D

OpenStudy (escamer):

@KendrickLamar2014

OpenStudy (escamer):

sorry :( i dont know this but kendrick will

OpenStudy (elisaneedshelp):

its okay :)

OpenStudy (elisaneedshelp):

@sleepyjess

OpenStudy (elisaneedshelp):

@IrishBoy123 can you see what is wrong :(

OpenStudy (irishboy123):

Three consecutive ***odd*** integers

OpenStudy (irishboy123):

they are 2 apart on the number line

OpenStudy (elisaneedshelp):

Im so confused tho. I didnt even know how to do it myself.

OpenStudy (irishboy123):

soz, you've lost me now if they are odd, they are 2 apart on the number line so it is n + (n + 2) + ((n + 2) + 2) = .......

OpenStudy (elisaneedshelp):

lol i guess ill leave this topic and do another one because I cant

OpenStudy (irishboy123):

well, do come back when you are ready i though you could already do the *consecutive* number problem the *consecutive* **odd** number problem is just an extension

OpenStudy (elisaneedshelp):

@ZacharySanders please dont stand by people's questions for no reason Its VERY irritating for like idk how many hours now :/